Assuming no need for any other nutrients, it has been calculated that an average human could meet their energy needs by eating 3 moles (540 g) of glucose. The complete oxidation of each molecule of glucose will result in the production of 30 molecules of ATP. The concentration of ATP within our cells remains constant at 2mM. Assuming an average adult has 25 litres of intracellular fluid calculate how many times per day each ATP molecule is hydrolysed and resynthesised.
Answer #1

1 molecule of glucose produce 30 molecules of ATP So, 3 moles of glucose produce ----> 3 x 30 = 90 moles of ATP.

For 2mM concernatation of ATP in 25 liter = 2× 10-3 × 25 = 0.05 moles of ATP.

So 90 -0.05 = 89.95 moles of ATP are hydrolysis and resynthesis in each day.

Or 89.95 × 6.022 × 1023 = 5.417 × 1025 molecules of ATP are gydrolhydr and resynthesis in each day.